Calculate the change in the entropy of the nitrogen during the transition below. A sample of nitrogen gas is heated from 20 centigrades to 400 centigrades at constant volume. Initial volume and pressure are given as 20liters and 5 kPa.The molar heat capacity of nitrogen at constant volume is 20.81 mole per kelvin per joule.


1
Expert's answer
2021-06-30T03:57:16-0400

Change in Entropy

=400-20 = 380

= 380 × 5 - 20.81 ×20

= 1900 - 416.2

= 1483.8 j/mol